# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ 3.1 Notations
Consider the sample space denoted as $\mathscr{T}$, which is a product of two distinct feature views, $A$ and $B$, expressed as $\mathscr{T}=\mathscr{T}^{A}\times\mathscr{T}^{B}$, where $\mathscr{T}^{A}\subseteq\mathbb{R}^{m_{1}}$, $\mathscr{T}^{B}\subseteq\mathbb{R}^{m_{2}}$ and $\mathscr{Y}=\{-1,+1\}$ denotes the label space. Here $n$ represents the number of samples, $m_{1}$ and $m_{2}$ denote the number of features corresponding to view $A$ (Vw-$A$) and view $B$ (Vw-$B$), respectively. Suppose =$\mathscr{H}=\{(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathscr{T}^{A},x_{i}^{B}\in\mathscr{T}^{B},y_{i}\in\mathscr{Y}\}_{i=1}^{n}$ represent a two-view dataset. Let $X_{1}\subseteq\mathbb{R}^{n\times m_{1}}$ and $X_{2}\subseteq\mathbb{R}^{n\times m_{2}}$ be the input matrix of Vw-$A$ and Vw-$B$, respectively and one-hot encoding matrix of the labels is denoted by $Y\subseteq\mathbb{R}^{n\times 2}$. $Z_{1}\subseteq\mathbb{R}^{n\times h_{l}}$ and $Z_{2}\subseteq\mathbb{R}^{n\times h_{l}}$ are hidden layer matrices, which are obtained by applying a nonlinear activation function, denoted as $\phi$, to the input matrices $X_{1}$ and $X_{2}$ after transforming them with randomly initialized weights and biases. Here $h_{l}$ represents the number of hidden layer nodes and (.$(.)^{t}$ represents the transpose operator.
3.2 Random Vector Functional Link (RVFL) Network
The RVFL (Pao et al., 1994) is a single-layer feed-forward neural network composed of three layers: the input layer, the hidden layer, and the output layer. The weights connecting the input and hidden layers, as well as the biases in the hidden layer, are randomly initialized and remain unchanged during the training process. The original features of input samples are also directly connected to the output layer. The determination of output layer weights involves analytical techniques such as the least square method or the Moore-Penrose inverse. The architecture of the RVFL model is shown in Fig. 1.
Figure 1: Geometrical structure of RVFL model
Consider $T=\{(x_{i},y_{i}),i=1,2,\ldots,n\}$ be the training dataset, where $y_{i}\in\{+1,-1\}$ represents the label of $x_{i}\in\mathbb{R}^{1\times m}$. Let $X=(x_{1}^{t},x_{2}^{t},\ldots,x_{n}^{t})^{t}\in\mathbb{R}^{n\times m}$ and $Y=(y_{1}^{t},y_{2}^{t},\ldots,y_{n}^{t})^{t}\in\mathbb{R}^{n\times 2}$ be the collection of all input and target vectors, respectively. $H_{1}$ is the hidden layer matrix, acquired by projecting the input matrix using randomly initialized weights and biases and then subjecting it to the non-linear activation function $\phi$. It is defined as:
$H_{1}=\phi(XW_{1}+b_{1})\in\mathbb{R}^{n\times h_{l}},$ (1)
where $W_{1}\in\mathbb{R}^{m\times h_{l}}$ represents the weights vector which is initialized randomly, and drawn from a uniform distribution spanning $[-1,1]$, and $b_{1}\in\mathbb{R}^{n\times h_{l}}$ is the bias matrix. Thus, $H_{1}$ is given as:
$\begin{bmatrix}X&H_{1}\end{bmatrix}W_{2}=\hat{Y}.$ (3)
Here, $W_{2}\in\mathbb{R}^{(m+h_{l})\times 2}$ denotes the weights matrix, which connects the input with the concatenation of the hidden nodes to the output nodes, while $\hat{Y}$ represents the predicted output. The resulting optimization problem of Eq. (3) is formulated as:
$(W_{2})_{min}=\left\{\begin{array}[]{ll}\left({H_{2}}^{t}{H_{2}}+\frac{1}{\mathcal{C}}I\right)^{-1}{H_{2}}^{t}{Y},&(m+h_{l})\leq n,\vspace{3mm}\\ {H_{2}}^{t}\left({H_{2}}{H_{2}}^{t}+\frac{1}{\mathcal{C}}I\right)^{-1}{Y},&n<(m+h_{l}),\end{array}\right.$ (7)
@@ 4 Data Preprocessing and Feature Extractions
In this section, we begin by outlining the DBP sequence from three distinct perspectives. Next, we utilize five different algorithms to extract features from these perspectives (Guan et al., 2022). Finally, we use MvRVFL to combine these features and develop the predictor for identifying DBPs. Figure 2 illustrates the flowchart of the model construction process.
@@ 4.1 Features of DBP
In this subsection, we detail the DBP sequence from three different perspectives: physicochemical properties, evolutionary information, and amino acid composition. These views are transformed into feature matrices using various extraction algorithms: Multi-scale Continuous and Discontinuous (MCD) (You et al., 2014) for amino acid composition; Normalized Moreau-Broto Autocorrelation (NMBAC) (Ding et al., 2016) for physicochemical properties; and Pseudo Position-Specific Scoring Matrix (PsePSSM) (Liu et al., 2015a), PSSM-based Discrete Wavelet Transform (PSSM-DWT) (Nanni et al., 2012), and PSSM-based Average Blocks (PSSM-AB) (cheol Jeong et al., 2010), for evolutionary information.
@@ 4.2 Multi-Scale Continuous and Discontinuous for Protein Sequences
Multi-scale Continuous and Discontinuous (MCD) uses multi-scale decomposition techniques to first segment the protein sequence into equal-length sections, followed by the characterization of each continuous and discontinuous region. Three descriptors are used for each region: distribution (D), composition (C), and transition (T). The following outlines the detailed process of feature extraction for each descriptor.
For distribution (D), the protein sequence is categorized into 7 different types of amino acids. The total count of amino acid occurrences in each category is represented by $M_{i},\leavevmode\nobreak\ i\in\{1,2,\ldots,7\}$. Subsequently, within each amino acid category, we determine the positions of the $1^{st}$, $25^{th}$ percentile, $50^{th}$ percentile, $75^{th}$ percentile, and last amino acids of that category in the entire protein sequence. These positions are then normalized by dividing by the total sequence length $L$. Therefore, each amino acid category can be represented by a $5$-dimensional feature vector. For the entire protein sequence, this results in a $35$-dimensional feature vector.
For composition (C), the 20 standard amino acids are grouped into seven categories based on the dipoles and volumes of their side chains, with each group sharing similar characteristics. Table 1 provides a detailed list of the amino acids in each category. Let $S=\{s_{1},s_{2},\ldots,s_{L}\}$ denotes a protein sequence, where $s_{i}$ represents the $i^{th}$ residue and $L$ is the total length of the sequence. Based on Table 1, we can represent $S$ as $S=\{s_{1},s_{2},\ldots,s_{L}\}$, where each residue $s_{i}$ is classified into one of the categories $\{1,2,3,\ldots,7\}$. We then calculate the proportion of residues in each category throughout the entire protein sequence. For composition (C), it is represented as a 7-dimensional vector, where each dimension corresponds to one of the seven amino acid categories.
@@
For transition (T), the entire protein sequence is divided into seven groups of amino acids according to the previously defined classification. We can calculate the frequency of transitions between different amino acid categories, which will be used to characterize the transition (T) properties of the protein. For transition (T), it can be represented as a $21$-dimensional vector, accounting for all possible transformations between the seven amino acid categories.
In this study, each protein sequence is partitioned into $5$ uniform-length segments, creating a total of 14 continuous or discontinuous regions. Each region is characterized by the $63$-dimensional features mentioned earlier ($7$ for composition, $21$ for transition, and $35$ for distribution). Consequently, the entire protein sequence is described using an $882$-dimensional feature vector ($63\times 14$).
Table 2: Six Physicochemical Properties Associated with Each Amino Acid
@@ 4.3 Normalized Moreau-Broto Autocorrelation for Protein Sequences
Normalized Moreau-Broto Autocorrelation (NMBAC) is a feature based on the physicochemical properties of amino acids. We use the same approach (Ding et al., 2016) to extract features from protein sequences. In this study, six physicochemical properties of amino acids are taken into account: hydrophobicity (H), solvent accessible surface area (SASA), polarizability (P1), polarity (P2), volume of side chains (VSC), and net charge index of side chains (NCISC). The values for these six physicochemical properties for each amino acid are listed in Table 2. We normalize these values for each physicochemical property using the following formula:
@@
where $M_{i,j}$ represents the value of the $i^{th}$ amino acid for physicochemical property $j$, and $M_{j}$ is the mean value of the $20$ amino acids for property $j$. $S_{j}$ is the standard deviation of the physicochemical property $j$ among the $20$ amino acids. Hence, NMBAC can be computed using the following formula:
@@
where $j\in\{1,2,\ldots,6\}$ represents one of the six physicochemical properties, $i\in\{1,2,\ldots,20\}$ denotes the $i^{th}$ amino acid in the protein sequence, $L$ is the length of the protein sequence, and $lg$ is the distance between residues. We use the same $lg$ values as those in (Ding et al., 2016), ranging from $1$ to $30$. Consequently, this results in a $180$-dimensional vector $(30\times 6)$. Furthermore, we incorporate the frequency of the $20$ amino acids found in the protein sequence into the NMBAC feature. As a result, the final NMBAC feature vector is $200$-dimensional $(180+20)$.
@@ 4.4 PSSM-Based Average Blocks for Protein Sequences
As PSSM includes evolutionary information about proteins, it has become a popular tool for various protein function prediction tasks in recent years. In this study, we extracted three features from PSSM: Pseudo Position-Specific Scoring Matrix (PsePSSM), PSSM-based Discrete Wavelet Transform (PSSM-DWT), and PSSM-based Average Blocks (PSSM-AB). The original PSSM profile, denoted as $Q_{PSSM}$, is represented as:
@@
The PSSM is initially divided into segments by PSSM-AB, where each block covers $5\%$ of the protein sequence. Consequently, the PSSM is divided into $20$ fundamental blocks, each of which consists of $20$ columns, independent of the length of the protein. The formula for extraction is as follows:
@@
where $E_{j}$ represents the feature vector of the $j^{th}$ block, with a dimension of $1\times 20$. The sequence length of the $j^{th}$ block is represented by $A_{j}$, and the PSSM value for the $i^{th}$ residue in the $j^{th}$ block, which also has a dimension of $1\times 20$, is indicated by $Q_{i}^{(j)}$. The feature vector derived from each protein sequence has a dimension of $400$ $(20\times 20)$, with a total of $20$ blocks.
@@ 4.5 PSSM-Based Discrete Wavelet Transform for Protein Sequences
We apply the Discrete Wavelet Transform (DWT) to extract features from the PSSM, referred to as PSSM-based DWT (PSSM-DWT). In this feature extraction method, each column of the PSSM profile of the protein serves as the input signal. Following the approach in (Ding et al., 2016), we apply a $4$-level discrete wavelet transform for analyzing the PSSM, as illustrated in Fig 3.
We use high-pass and low-pass filters to break down the approximation function at each step of the process, followed by downsampling. For every level, this method yields low-frequency coefficients $L_{i}$ and high-frequency coefficients $H_{i}$, where $i$ denotes the $i^{th}$ layer. We calculate the mean, median, maximum, and minimum values for both the low-frequency and high-frequency coefficients at each layer. Furthermore, we extract additional specific information from each layer of the low-frequency component by extracting the first five discrete cosine coefficients. Thus, for a protein’s PSSM profile, we obtain a $1040$-dimensional feature vector $((4+4+5)\times 4\times 20)$.
@@ 4.6 Pseudo PSSM for Protein Sequences
$Q_{PsePSSM}=[Q_{1}\leavevmode\nobreak\ Q_{2}\leavevmode\nobreak\ \ldots Q_{20}\leavevmode\nobreak\ J_{1}^{\zeta_{1}}\leavevmode\nobreak\ J_{2}^{\zeta_{1}}\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ J_{20}^{\zeta_{1}}\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ J_{1}^{\zeta_{N}}\leavevmode\nobreak\ J_{2}^{\zeta_{N}}\leavevmode\nobreak\ \ldots\leavevmode\nobreak\ J_{20}^{\zeta_{N}}]^{T},$ (12)
@@
$Q_{j}=\frac{1}{L}\sum_{i=1}^{L}Q_{i,j},\leavevmode\nobreak\ \leavevmode\nobreak\ j=1,2,\ldots,20,$ (13)
@@
[…]
where $L$ denotes the length of the protein sequence, $l$ represents the distance between residues, $Q_{j}$ indicates the average score of residues in the protein sequence that evolve into the $j^{th}$ residue, and $J_{j}^{\zeta_{l}}$ reflects the PSSM scores for two residues that are adjacent with a distance of $l$. We adopt the same range of $l$ values as those used in (Chou & Shen, 2007), where $l$ varies from 1 to $15$. Thus, for a protein sequence, we generate a $320$-dimensional vector $(20+20\times 15)$. We can extract PSSM features with a fixed length using this technique while preserving sequence order information.
@@ 4.7 Feature Selection for Protein Sequences
In this section, we discuss feature selection and our use of the Minimum Redundancy–Maximum Relevancy (mRMR) method (Wei et al., 2017). This approach ranks the importance of input features by minimizing redundancy among them and maximizing their relevance to the target. In mRMR, mutual information is utilized to compute the redundancy and relevance mentioned above. The calculations are performed as follows:
$J(g,h)=\iint p(g,h)\log\frac{p(g,h)}{p(g)p(h)}dg\leavevmode\nobreak\ dh,$ (15)
where $g$ and $h$ are two vectors, $p(g)$ and $p(h)$ represent the marginal probability densities of $g$ and $h$ respectively, and $p(g,h)$ denotes the joint probability density of $g$ and $h$. We use $T$ to denote the full set of features, $T_{t}$ to represent a sorted subset with $m$ features, and $T_{n}$ to indicate an unsorted subset with $n$ features. Thus, the relevance $L$ and redundancy $D$ of features $k$ within the subset $T_{n}$ are calculated as follows:
@@
$\underset{k_{i}\in T_{n}}{\min}[D(k_{i})-L(k_{i})],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,3,\ldots,n.$ (18)
By applying mRMR, we generate a reordered set of features, with each feature ranked according to its importance. For each feature, we subsequently select the optimal subset to be used in further experiments. Among the five feature types used in this paper, we first ranked the importance of features within each type using mRMR. Among the five feature types examined in this paper, we first ranked the importance of features within each category using mRMR. We then selected the top $1/2$, $3/4$, $7/8$, and $15/16$ of the features from the ranked list to form new feature subsets, respectively.

### v2

@@ 2 Notations
Let $\mathscr{T}=\mathscr{T}^{A}\times\mathscr{T}^{B}$ be the sample space defined by the Cartesian product of two separate feature views, $A$ and $B$, where $\mathscr{T}^{A}\subseteq\mathbb{R}^{m_{1}}$, $\mathscr{T}^{B}\subseteq\mathbb{R}^{m_{2}}$ and $\mathscr{Y}=\{-1,+1\}$ denotes the label space. Here $n$ represents the number of samples, $m_{1}$ and $m_{2}$ denote the number of features corresponding to view AA (Vw-AA) and view $B$ (Vw-$B$), respectively. Suppose $\mathscr{H}=\{(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathscr{T}^{A},x_{i}^{B}\in\mathscr{T}^{B},y_{i}\in\mathscr{Y}\}_{i=1}^{n}$ represent a two-view dataset. Let $X_{1}\subseteq\mathbb{R}^{n\times m_{1}}$ and $X_{2}\subseteq\mathbb{R}^{n\times m_{2}}$ be the input matrix of Vw-$A$ and Vw-$B$, respectively, and $Y$ denotes the label vector. $H_{1}\subseteq\mathbb{R}^{n\times h_{l}}$ and $H_{2}\subseteq\mathbb{R}^{n\times h_{l}}$ are hidden layer matrices, which are obtained by applying a nonlinear activation function, denoted as $\phi$, to the input matrices $X_{1}$ and $X_{2}$ after transforming them with randomly initialized weights and biases. Here, $h_{l}$ denotes the number of nodes in the hidden layer, and $(\cdot)^{t}$ indicates the transpose operation.
@@ 3 Data Preprocessing and Feature Extractions
In this part, we first describe the DBP sequence from three different viewpoints. Next, five distinct algorithms are employed to derive features from these various viewpoints [25]. Finally, the MvRVFL model is leveraged to integrate the extracted features and construct the predictor for detecting DNA-binding proteins (DBPs). Fig 1 illustrates the flowchart of the model construction process.
@@ 3.1 Features of DBP
In this subsection, we detail the DBP sequence from three different perspectives: physicochemical properties, evolutionary information, and amino acid composition. These views are transformed into feature matrices using various extraction algorithms: Multi-scale Continuous and Discontinuous (MCD) [73] for amino acid composition; Normalized Moreau-Broto Autocorrelation (NMBAC) [14] for physicochemical properties; and PSSM-based Discrete Wavelet Transform (PSSM-DWT) [44], Pseudo Position-Specific Scoring Matrix (PsePSSM) [37], and PSSM-based Average Blocks (PSSM-AB) [7], for evolutionary information.
@@ 3.2 Multi-Scale Continuous and Discontinuous for Protein Sequences
The Multi-scale Continuous and Discontinuous (MCD) approach employs multi-scale decomposition to partition a protein sequence into equal-length segments, subsequently analyzing both continuous and discontinuous regions within each segment. Three descriptors are used for each region: distribution (D), composition (C), and transition (T). The following outlines the detailed process of feature extraction for each descriptor.
For distribution (D), the protein sequence is categorized into 7 different types of amino acids. The total number of occurrences of amino acids in each category is denoted by $M_{i}$, where $i$ belongs to the set $\{1,2,\ldots,7\}$. Subsequently, within each amino acid category, we determine the positions of the $1^{st}$, $25^{th}$ percentile, $50^{th}$ percentile, $75^{th}$ percentile, and last amino acids of that category in the entire protein sequence. These positions are then normalized by dividing by the total sequence length $L$. Therefore, each amino acid category can be represented by a $5$-dimensional feature vector. For the entire protein sequence, this results in a $35$-dimensional feature vector.
In the Composition (C) approach, the 20 standard amino acids are classified into seven groups according to their dipole moments and side-chain volumes, with each group exhibiting comparable characteristics. Table 1 outlines the specific amino acids assigned to each respective group in detail. Let $S=\{s_{1},s_{2},\ldots,s_{L}\}$ represent a protein sequence, where each $s_{i}$ corresponds to the $i^{\text{th}}$ amino acid residue, and $L$ indicates the overall sequence length. Based on Table 1, we can represent $S$ as $S=\{s_{1},s_{2},\ldots,s_{L}\}$, where each residue $s_{i}$ is classified into one of the categories $\{1,2,3,\ldots,7\}$. We calculate the proportion of amino acid residues in each category throughout the entire length of the protein sequence. In the case of composition (C), it is expressed as a 7-dimensional feature vector, with each component representing one of the seven predefined amino acid groups.
@@
For transition (T), the protein sequence is segmented into seven groups of amino acids according to the previously defined classification. The transition frequency between various amino acid categories can be computed to represent the protein’s transition (T) characteristics. This transition information is captured using a 21-dimensional vector, which encompasses all possible pairwise transitions among the seven amino acid groups.
In this study, each protein sequence is partitioned into $5$ uniform-length segments, creating a total of 14 continuous or discontinuous regions. Each region is characterized by the $63$-dimensional features mentioned earlier ($35$ for distribution, $21$ for transition, and $7$ for composition). Consequently, the entire protein sequence is described using an $882$-dimensional feature vector ($63\times 14$).
Table 2: Six physicochemical characteristics linked to each amino acid.
@@ 3.3 Normalized Moreau-Broto Autocorrelation for Protein Sequences
The Normalized Moreau-Broto Autocorrelation (NMBAC) is a descriptor derived from the physicochemical attributes of amino acids. Following the methodology outlined in [14], we extract relevant features from protein sequences. Specifically, this work considers six physicochemical characteristics: hydrophobicity (O), solvent-accessible surface area (SASA), polarizability (P1), polarity (P2), side chain volume (VSC), and the net charge index of side chains (NCISC). Table 2 lists the values of these six physicochemical properties for each amino acid. The values for each property are standardized by the given equation:
@@
where $M_{i,j}$ denotes the value assigned to the $i^{th}$ amino acid corresponding to the $j^{th}$ physicochemical property, while $M_{j}$ refers to the average of that property calculated across all 20 standard amino acids. $S_{j}$ represents the standard deviation of property $j$ among these 20 amino acids. Hence, NMBAC can be computed using the following formula:
@@
where $j\in\{1,2,\ldots,6\}$ corresponds to one of the six selected physicochemical properties, $i\in\{1,2,\ldots,20\}$ indicates the $i^{th}$ amino acid in the sequence, $L$ stands for the total length of the protein sequence, and $lg$ refers to the lag or interval between amino acid residues. We use the same $lg$ values as those in [14], ranging from $1$ to $30$. Consequently, this results in a $180$-dimensional vector $(30\times 6)$. Moreover, the NMBAC feature incorporates the occurrence frequencies of the 20 amino acids in the protein sequence. As a result, the NMBAC feature vector has a total dimension of $200(180+20)$.
@@ 3.4 PSSM-Based Average Blocks for Protein Sequences
Due to its ability to encapsulate evolutionary insights, the Position-Specific Scoring Matrix (PSSM) has gained widespread use in recent years for predicting protein functions. Three different feature types are taken from the PSSM: Pseudo-PSSM (PsePSSM), PSSM-derived Average Blocks (PSSM-AB), and Discrete Wavelet Transform based on PSSM (PSSM-DWT). The original PSSM matrix, denoted as $Q_{PSSM}$, is defined as follows:
@@
The PSSM is first partitioned into segments by PSSM-AB, with each segment representing $5\%$ of the total protein sequence. As a result, the PSSM is split into $20$ core blocks, each containing $20$ columns, regardless of the protein’s length. The extraction formula is given by:
@@
where $E_{j}$ represents the feature vector of the $j^{th}$ block, and its dimension is $1\times 20$. The sequence length in the $j^{th}$ block is denoted by $A_{j}$, and $Q_{i}^{(j)}$ refers to the PSSM value for the $i^{th}$ residue within the $j^{th}$ block, also with a dimension of $1\times 20$. The feature vector derived from each protein sequence has a dimension of $400$ $(20\times 20)$, with a total of $20$ blocks.
@@ 3.5 PSSM-Based Discrete Wavelet Transform for Protein Sequences
To derive features using the Discrete Wavelet Transform (DWT), we treat each column of the protein’s PSSM profile as an input signal. This approach, known as PSSM-based DWT (PSSM-DWT), enables the extraction of informative patterns from the PSSM. As shown in Fig. 2, a 4-level discrete wavelet transform is applied to the PSSM for analysis [14].
High-pass and low-pass filters are used to partition the approximation function at each stage of the procedure, followed by downsampling. As a result, high-frequency coefficients $O_{i}$ and low-frequency coefficients $L_{i}$ are produced, where $i$ indicates the decomposition level. For each set of coefficients at every layer, we compute statistical measures including the mean, median, maximum, and minimum. Furthermore, we extract more information from each layer of the low-frequency component by obtaining the first five discrete cosine transform coefficients. Thus, for a protein’s PSSM profile, we obtain a $1040$-dimensional feature vector $((4+4+5)\times 4\times 20)$.
@@ 3.6 Pseudo PSSM for Protein Sequences
$Q_{PsePSSM}=[Q_{1}\penalty 10000\ Q_{2}\penalty 10000\ \ldots Q_{20}\penalty 10000\ J_{1}^{\zeta_{1}}\penalty 10000\ J_{2}^{\zeta_{1}}\penalty 10000\ \ldots\penalty 10000\ J_{20}^{\zeta_{1}}\penalty 10000\ \ldots\penalty 10000\ J_{1}^{\zeta_{N}}\penalty 10000\ J_{2}^{\zeta_{N}}\penalty 10000\ \ldots\penalty 10000\ J_{20}^{\zeta_{N}}]^{T},$ (5)
@@
$Q_{j}=\frac{1}{L}\sum_{i=1}^{L}Q_{i,j},\penalty 10000\ \penalty 10000\ j=1,2,\ldots,20,$ (6)
@@
$J_{j}^{\zeta_{l}}=\frac{1}{L-\zeta_{l}}\sum_{i=1}^{L-\zeta_{l}}[Q_{i,j}-Q_{(i+\zeta_{l}),j}]^{2},\penalty 10000\ \penalty 10000\ \penalty 10000\ \zeta_{l}<L;\penalty 10000\ 1\leq l\leq N,$ (7)
where $l$ indicates the gap between residues, $L$ denotes the total length of the protein sequence, and $Q_{j}$ represents the average score for residues in the sequence transitioning to the $j^{th}$ residue, and $J_{j}^{\zeta_{l}}$ denotes the PSSM values for two adjacent residues separated by a distance of $l$. We adopt the same range of $l$ values as those used in [10], where $l$ varies from 1 to $15$. As a result, we form a vector of size 320 for each protein sequence, which consists of $20$ plus $20\times 15$ components. We can extract PSSM features with a fixed length using this technique while preserving sequence order information.
@@ 3.7 Feature Selection for Protein Sequences
In this section, we discuss feature selection and our use of the Minimum Redundancy–Maximum Relevancy (mRMR) method [67]. This method prioritizes input features by reducing redundancy while enhancing their relevance to the target. In mRMR, mutual information helps evaluate the relevance and redundancy of the features. The calculations are performed as follows:
$J(g,h)=\iint p(g,h)\log\frac{p(g,h)}{p(g)p(h)}dg\penalty 10000\ dh,$ (8)
where $g$ and $h$ be two vectors, with $p(g)$ and $p(h)$ denoting their respective marginal probability distributions, and $p(g,h)$ indicating their joint probability distribution. The complete feature set is represented by $T$, while $T_{t}$ refers to an ordered subset containing $m$ features, and $T_{n}$ denotes an unordered subset comprising $n$ features. Therefore, the relevance $L$ and redundancy $D$ of features $k$ in the subset $T_{n}$ are determined using the following formulas:
@@
$\underset{k_{i}\in T_{n}}{\min}[D(k_{i})-L(k_{i})],\penalty 10000\ \penalty 10000\ \penalty 10000\ i=1,2,3,\ldots,n.$ (11)
Using mRMR, we create a reorganized feature set, with each feature ranked based on its significance. For each type of feature considered in this study, we began by evaluating the importance of individual features using the mRMR algorithm. Based on the resulting ranking, we constructed several new subsets by selecting the top-ranked features, specifically the top $1/2$, $3/4$, $7/8$, and $15/16$ portions of the full ranked feature list, which were then used for further experimentation.